# CoLLEGe: Concept Embedding Generation for Large Language Models

### v1

@@ Acknowledgment
Appendix A GRE Verbal Reasoning Examples
The GRE task consists of two different types of fill-in-the-blank questions. The first type asks you to select the best possible choice to complete the provided sentence, so it serves as a test of the top-1 prediction. The second type asks for which two words best complete the sentence. Often these words are similar, but distinct. It tests how the top-2 predictions of the LLM.
@@ Appendix A GRE Verbal Reasoning Examples
Mary’s former classmates were taken aback by her [BLANK] behavior at the reunion for, during her school years, she was frequently reprimanded for creating disturbances with her exuberant outbursts and playful antics.
a) gregarious b) discourteous c) obsequious d) reticent e) scurrilous
d) reticent
Choose the word for each blank that best fits the meaning of the sentence as a whole.
The firefighter, desperate to save the children on the second floor of the fiery house, rushed into their bedroom; his colleagues, more wary of the [BLANK] structure, remained outside.
a) stalwart b) precarious c) stout d) irrefragable e) tottering f) fecund
b) precarious e) tottering
Select the two answer choices that, when inserted into the sentence, fit the meaning of the sentence as a whole and yield complete sentences that are similar in meaning.
Table 7: Examples of both types of questions for the GRE task dataset
Appendix B Data Processing
@@
Appendix C Baseline Implementations
@@
For Token Tuning, I treat the example sentences as a ”batch” and perform N=1,2 steps of gradient descent on only the input and output embeddings for the new token.
Appendix D Results for Prompting+
@@
D.1 GRE Results for Prompting+
We show results for the GRE task when prompting with examples in-context, using embeddings for the new token generated by the model or baseline following the “+”.
@@ D.1 GRE Results for Prompting+
Method 1-Shot 2-Shot 3-Shot 4-Shot Def + 1-Shot Def + 2-Shot Def + 3-Shot Def + 4-Shot
@@
All CoLLEGe models outperform the Prompting baseline, where new token embeddings are randomly initialized. Each “Prompting+CoLLEGe” model performs worse than the unprompted version, which we hypothesize is due to a lack of instruction tuning data in the training dataset.
D.2 Definition Generation Results for Prompting+
We present results for our definition generation task with examples presented in-context, using our baselines or CoLLEGe to generate the new token embedding. ELO scores are calculated separately from those in Table 4 by sampling a random baseline challenger to the “Prompting+CoLLEGe” model.
@@ D.2 Definition Generation Results for Prompting+
Prompting 0.825 $\pm$ 0.028 1002.09
$\pm$
22.87
 + TT-1 0.762 $\pm$ 0.035 959.88
$\pm$
17.22
 + TT-2 0.763 $\pm$ 0.045 962.10
$\pm$
12.65
 + HiCE 0.740 $\pm$ 0.098 949.32
$\pm$
10.77
 + Additive 0.735 $\pm$ 0.098 950.61
$\pm$
3.48
 + CoLLEGe 0.858 $\pm$ 0.029 1176.00
$\pm$
12.69
@@
D.3 Twitter Results for Prompting+
When examples Tweets are provided in-context, our CoLLEGe model’s accuracy on the slang identification task increases. For other baselines, aside from prompting with randomly initialized embeddings, however, performance either degrades or remains about the same. With the Word2Vec-based baselines, this may be due to the difficulty of mapping between Word2Vec embeddings and the LLaMA input and output embedding space.
@@ D.3 Twitter Results for Prompting+
Appendix E Generated Definitions
E.1 Additional CoLLEGe Definitions
We show additional definitions generated from CoLLEGe, including definitions generated with more than one example. CoLLEGe is able to generate plausible definitions that capture important features of the new word. In some cases, the CoLLEGe definition is incorrect, but plausible based on the examples provided. For example, the generated definition for opera company is incorrect, but reasonable in light of the example sentences.
@@ E.1 Additional CoLLEGe Definitions
During the complex abdominal surgery, the surgeon carefully moved the <nonce> aside to gain better access to the patient’s damaged organs.
a surgical procedure in which a portion of the intestine is brought through an opening in the abdominal wall.
a fold of peritoneum supporting the viscera
omentum
The yellow blooms of the <nonce> added a vibrant contrast to the green canvas of the wetlands.
”a plant of the genus Ficus, having a milky sap and large, often edible, fruit.”
aromatic evergreen or deciduous dioecious shrubs or trees of eastern Asia and North America
Lindera
@@
a person who is skilled in the art of dancing.
a company that produces operas
opera
@@
the most beautiful or perfect specimen of its kind.
informal usage attributing authenticity
real McCoy
Table 11: Additional definitions generated with CoLLEGe, using the prompt “The word |<nonce>| is defined as”. Each definition is generated using the examples. None of the example sentences are provided in-context to the model.
E.2 Qualitative Comparison of Generated Definitions
We present a side-by-side comparison of definitions generated by CoLLEGe as well as the baselines in Table 12. Definitions generated by the baselines (with the exception of Prompting) are essentially unusable.
Word/
Phrase
@@ E.2 Qualitative Comparison of Generated Definitions
CoLLEGe HiCE Additive TT-1 TT-2
horsecar
an early form of streetcar that was drawn by horses
a “motorized vehicle with a cabin and a platform for passengers to stand on.”
a word” “the a word a” is a
that the place where you are the place.
follows.
follows.
popishly
like the Pope; in a popish manner
in a manner that is intended to attract attention or admiration.
a ” a “a” a a word a word a a a
that the same.
’the opposite of the word ” ” in the dictionary
the opposite of the word,
Table 12: Side-by-side comparison between CoLLEGe-generated definitions and definitions generated by the baselines without prompting.
E.3 Failures of Generated Definitions
@@
Appendix F Slang Sources
To build the dataset of slang terms and definitions, we used UrbanDictionary [^5] https://www.urbandictionary.com/ Dictionary.com’s Pop Culture [^6] https://www.dictionary.com/e/pop-culture/ and Slang [^7] https://www.dictionary.com/e/slang/ sections, the recent American Dialect Society meeting [^8] https://americandialect.org/nominations-for-words-of-the-year-2023/, Bark 999bark.us, Wiktionary [^10] https://www.wiktionary.org/, and the Online Slang Dictionary [^11] http://onlineslangdictionary.com/

### v2

@@ Acknowledgments
Appendix A Dataset Composition
Our dataset for training CoLLEGe consists of samples drawn from subsets of the Pile shown below in Table 6.
Source Num. Examples
Pile-CC 79,606
Books3 51,850
BooksCorpus2 3,436
Table 6: The top Pile subsets represented in our dataset.
Appendix B Implementation Details
During training, we provide 1 support sequence for each query sequence and generalize to $K>1$ during testing. We train our model for 28000 steps at batch size 32, with a learning rate of 1e-3, a linear learning rate schedule with warmup, and using the AdamW optimizer (Loshchilov & Hutter, 2019) with weight decay = 0.1 and default beta values. We experimented with different beta values during training, but found they had little effect. During training we clip gradients to a norm of 1.0. The final model checkpoint is selected based on its GRE score.
Appendix C GRE Verbal Reasoning Examples
The GRE task consists of two different types of fill-in-the-blank questions. The first type asks you to select the best possible choice to complete the provided sentence, so it serves as a test of the top-1 prediction. The second type asks for which two words best complete the sentence. Often these words are similar, but distinct. It tests how the top-2 predictions of the LLM. We show an example for both types below in Table 7.
@@ Appendix C GRE Verbal Reasoning Examples
Mary’s former classmates were taken aback by her [BLANK] behavior at the reunion for, during her school years, she was frequently reprimanded for creating disturbances with her exuberant outbursts and playful antics. a) gregarious b) discourteous c) obsequious d) reticent e) scurrilous d) reticent Choose the word for each blank that best fits the meaning of the sentence as a whole.
The firefighter, desperate to save the children on the second floor of the fiery house, rushed into their bedroom; his colleagues, more wary of the [BLANK] structure, remained outside. a) stalwart b) precarious c) stout d) irrefragable e) tottering f) fecund b) precarious e) tottering Select the two answer choices that, when inserted into the sentence, fit the meaning of the sentence as a whole and yield complete sentences that are similar in meaning.
Table 7: The two types of questions used in CoLLEGe-GRE.
Appendix D Data Processing
@@
We used slang definitions from UrbanDictionary [^5] https://www.urbandictionary.com/ Dictionary.com’s Pop Culture [^6] https://www.dictionary.com/e/pop-culture/ and Slang [^7] https://www.dictionary.com/e/slang/ sections, the recent American Dialect Society meeting [^8] https://americandialect.org/nominations-for-words-of-the-year-2023/, Bark 999bark.us, Wiktionary [^10] https://www.wiktionary.org/, and the Online Slang Dictionary [^11] http://onlineslangdictionary.com/ for our dataset.
Appendix E Baseline Implementations
@@
For Token Tuning, we treat the example sentences as a ”batch” and perform N=1,2 steps of gradient descent on only the input and output embeddings for the new token.
Appendix F Results for Prompting+
@@
F.1 GRE Results for Prompting+
Table 8 shows results for the GRE task when prompting with examples in-context, using embeddings for the new token generated by the model or baseline following the “+”.
@@ F.1 GRE Results for Prompting+
Method 1-Shot 2-Shot 3-Shot 4-Shot D+1-Shot D+2-Shot D+3-Shot D+4-Shot
@@
All CoLLEGe models outperform the Prompting baseline, where new token embeddings are randomly initialized. Prompting by including the definition of each term alongside the example sentences improves performance for the baselines the most, but CoLLEGe significantly outperforms. Each “Prompting+CoLLEGe” model performs worse than the unprompted version, which we hypothesize is due to a lack of instruction tuning data in the training dataset.
F.2 Definition Generation Results for Prompting+
We present results in Table 9 for our definition generation task with examples presented in-context, using our baselines or CoLLEGe to generate the new token embedding. ELO scores are calculated separately from those in Table 2 by sampling a random baseline challenger to the “Prompting+CoLLEGe” model.
@@ F.2 Definition Generation Results for Prompting+
Prompting 0.825 $\pm$ 0.028 1002.09 $\pm$ 22.87
   + TT-1 0.762 $\pm$ 0.035 959.88 $\pm$ 17.22
   + TT-2 0.763 $\pm$ 0.045 962.10 $\pm$ 12.65
   + HiCE 0.740 $\pm$ 0.098 949.32 $\pm$ 10.77
   + Additive 0.735 $\pm$ 0.098 950.61 $\pm$ 3.48
   + CoLLEGe 0.858 $\pm$ 0.029 1176.00 $\pm$ 12.69
@@
F.3 Twitter Results for Prompting+
When examples Tweets are provided in-context, our CoLLEGe model’s accuracy on the slang identification task increases. For other baselines, aside from prompting with randomly initialized embeddings, however, performance either degrades or remains about the same. With the Word2Vec-based baselines, this may be due to the difficulty of mapping between Word2Vec embeddings and the LLaMA input and output embedding space. We present these results in Table 10.
@@ F.3 Twitter Results for Prompting+
Appendix G Generated Definitions
G.1 Additional CoLLEGe Definitions
We show additional definitions generated from CoLLEGe, including definitions generated with more than one example, in Table 11.
@@ G.1 Additional CoLLEGe Definitions
During the complex abdominal surgery, the surgeon carefully moved the <nonce> aside to gain better access to the patient’s damaged organs. a surgical procedure in which a portion of the intestine is brought through an opening in the abdominal wall. a fold of peritoneum supporting the viscera omentum
The yellow blooms of the <nonce> added a vibrant contrast to the green canvas of the wetlands. ”a plant of the genus Ficus, having a milky sap and large, often edible, fruit.” aromatic evergreen or deciduous dioecious shrubs or trees of eastern Asia and North America Lindera
@@
a person who is skilled in the art of dancing. a company that produces operas opera
@@
the most beautiful or perfect specimen of its kind. informal usage attributing authenticity real McCoy
Table 11: Additional definitions generated with CoLLEGe, using the prompt “The word <nonce> is defined as”. Each definition is generated using the examples. None of the example sentences are provided in-context to the model.
Word/Phrase horsecar popishly
@@
an early form of streetcar that was drawn by horses like the Pope; in a popish manner
CoLLEGe a “motorized vehicle with a cabin and a platform for passengers to stand on.” in a manner that is intended to attract attention or admiration
HiCE a word” “the a word a” is a a ” a “a” a a word a word a a a
Additive that the place where you are the place. that the same.
TT-1 follows. ’the opposite of the word ” ” in the dictionary
TT-2 follows. the opposite of the word,
Prompting in the heart of bustling 19th century Boston, he spent his mornings as a loyal conductor, navigating through a flamboyant manner of acting or speaking.
Table 12: Side-by-side comparison between CoLLEGe-generated definitions and definitions generated by the baselines.
G.2 Qualitative Comparison of Generated Definitions
We present the definitions generated by CoLLEGe side-by-side with those generated by our baselines in Table 12.
In general, the baselines (without Prompting) are unable to generate usable definitions for the new concept.
G.3 Failures of Generated Definitions
@@
Appendix H Slang Examples
We highlight a few Twitter slang examples from the hand-curated subset of our task dataset in Table 13.
Slang Term Definition Example Tweet
rizz The ability to confidently approach people and talk to them, in a more romantic or flirty way. Imagine having so little rizz that even the AI girlfriend rejects you. Just complete negative game.
hits different When something is significantly better than usual or is way better under certain circumstances. getting called pretty in person just hits different. people be making my day.
gorpcore A fashion style that is similar to that of
hiking/wilderness/utility tech wear.